please can somebody help with this I'm stuck
x-2/3 = 2x/7
if you can help please show how to work out
so I can understand

Answer by John10(297)

(x-2)/3 = 2x/7
Use the cross product:
7(x - 2) = 3(2x)
7x - 14 = 6x
7x - 6x = 14
x = 14
John10:)